A particule leaves the origin with a speed of 3×10^6m/s at 35° to the x-axis, it moves in a constant electric field E=E,j. Find Ey such that the particle will cross the x-axis at x=15cm if the particle is (a) an electron (b) a proton

1
Expert's answer
2021-11-24T11:58:26-0500

a)


"E=\frac{(3\cdot10^{6})^2\sin{70}}{(1.7588\cdot10^{11})(0.15)}=320\frac{V}{m}"

b)


"E=\frac{(3\cdot10^{6})^2\sin{70}}{(95756)(0.15)}=5.9\cdot10^{8}\frac{V}{m}"